OpenStudy (anonymous):

Can someone explain the One-to-One Property to me? And how I would use it to solve an equation for x? For example: 45) 3^x+1=27

OpenStudy (anonymous):

@DisplayError

OpenStudy (displayerror):

The one-to-one property says that if we're given a logarithmic function like this: \[5^{x-1} = 25\] And we can change the base of the right side to be the same as that on the left side, \[5^{x-1} = 5^2\] We can just set the exponents equal to each other to solve: \[x - 1 = 2\] \[x = 3\] The same can be done with logarithms. As long as the base is the same \[\log_5 8 = \log_5 (2x+2)\] We can set the argument of the logarithms to be equal in order to solve: \[8 = 2x+2\] \[x = 3\] Now for your given problem, we can apply the same idea \[3^x + 1 = 27\] Recall your exponents with base 3: \[\begin{array}{rcr} 3^0 & = & 1 \\ 3^1 & = & 3 \\ 3^2 & = & 9 \\ 3^3 & = & 27\end{array} \] With this information, you should be able to rewrite your equation such that they're all of the number 3 raised to an exponent. Then you should be able to solve for x using the one-to-one property.

OpenStudy (anonymous):

Ok.... that kind of explains it. Lol I'm still kinda confused. I think I might just ask my teacher to explain it to me tomorrow. Thanks anyways :)